# Does equivariance matter at scale?

### v1

@@ 4.2 Scaling with data
Next, we turn to the scaling with training data for a fixed training compute budget. In Fig. 2 we show the test loss as a function of the number of unnique training tokens. We compare baseline and equivariant transformers, each using a compute-optimal model size and training tokens for a training compute budget of $10^{18}$ nominal FLOPs.
@@
In Fig. 2 we also show results for a baseline transformer model trained with data augmentation. As expected, data augmentation does not make a difference when training for a single epoch. However, it drastically improves the performance in the small-data regime: when training for thousands of epochs, data augmentation makes a baseline transformer as data-efficient as an equivariant model.
@@ 5 Discussion
The first (and expected) benefit for the equivariant architecture is that it performs better than a non-equivariant architecture when only little training data is available, as we show in Fig. 2. However, we find a non-equivariant model trained with data augmentation performs just as well as the equivariant architecture, at least when the number of epochs (i. e. repeated uses of the same training sample) is sufficiently large.
@@
We hypothesize three possible explanations for this observation. First, the baseline transformer, the more mature architecture, may have a better initialization scheme and thus require less training steps to reach a good performance. Second, the different trade-offs may be related due to the different choice of width and depth between the architectures. A third possible explanation is linked to the internals of the equivariant transformer architecture, which can express certain primitives particularly efficiently: the free movement and gravitational acceleration of rigid bodies can be represented with few multivector channels, thanks to the geometric product operation integrated into the architecture. This explains why the architecture can achieve a good performance with very few parameters. However, lowering the loss further requires precise collision detection and modelling. These need substantially more computational operations and a substantial amount of scalar channels, similar to the non-equivariant transformer. This offers a possible explanation for why at a larger compute budget, a model size closer to that of the baseline transformer is compute-optimal.
@@
As much as we would like to, we cannot conclusively settle the question raised in the title of this paper. Our work is limited in several ways. First, we only analyzed a single benchmark problem and two model families. We chose a task with a common symmetry group and general-purpose architectures that are frequently applied to a wide range of problems. We believe it is important to study to what extent our findings generalize to other problems or to other architectures, for instance those based on message-passing over graphs. Moreover, on the problem we studied, we did not set a new state of the art: we deliberately focused on general-purpose models, which do not achieve the same level of performance as highly problem-specific architectures (Allen et al., 2022).
@@
Keeping these limitations in mind, we believe that our findings provide some evidence that symmetry-aware modelling can be a sensible choice even for large compute and data budgets. The benefits and disadvantages of strong inductive biases at scale are important for problems spanning several fields of science and engineering. We hope that our study can encourage further investigations into this question.
@@ Appendix A Dataset
We generate a benchmark dataset of rigid-body interactions with the Kubric simulator (Greff et al., 2022), which is based on the PyBullet physics engine (Coumans & Bai, 2016–2024). We follow the MOVi-B configuration as used by Allen et al. (2022): we generate trajectories trajectories of 96 frames at 48 frames per seconds. Our training set consists of $4\cdot 10^{5}$ such trajectories, while we use $1000$ trajectories each for the validation and test set.
@@ Appendix B Models
For the baseline transformer, these features are embedded using random Fourier features (Tancik et al., 2020) with 128 frequencies sampled from a Gaussian with standard deviation $0.1$.
@@
We also experimented with predicting all vertex positions directly, without enforcing object rigidiy, as well as with parametrizing elements of the Lie algebra of $\mathrm{E}(3)$, which would then be exponentiated to construct transformations $h$. Both approaches performed worse in initial tests.

### v2

@@ 4.2 Scaling with data
Next, we turn to the scaling with training data for a fixed training compute budget. In Fig. 2 we show the test loss as a function of the number of unique training tokens. We compare baseline and equivariant transformers, each using a compute-optimal model size and training tokens for a training compute budget of $10^{18}$ nominal FLOPs.
@@
In Fig. 2 we also show results for a baseline transformer model trained with data augmentation. As expected, data augmentation does not make a difference when training for a single epoch on a large dataset, which is the setting towards the right side of Fig. 2. However, data augmentation drastically improves the performance in the small-data regime: when training for thousands of epochs on a small dataset, data augmentation makes a baseline transformer as data-efficient as an equivariant model. This is the setting we see towards the left of Fig. 2.
@@ 5 Discussion
The first (and expected) benefit for the equivariant architecture is that it performs better than a non-equivariant architecture when only little training data is available, as we show in Fig. 2.
However, we find a non-equivariant model trained with data augmentation performs just as well as the equivariant architecture, at least when the number of epochs (i. e. repeated uses of the same training sample) is sufficiently large. Our findings suggest that non-equivariant architectures are indeed able to learn symmetries from data if they are shown sufficiently many symmetry-transformed examples.
@@
We hypothesize three possible explanations for this observation. First, the baseline transformer, the more mature architecture, may have a better initialization scheme and thus require less training steps to reach a good performance. Second, the different trade-offs may be related due to the different choice of width and depth between the architectures. A third possible explanation is linked to the internals of the equivariant transformer architecture, which can express certain primitives particularly efficiently: the free movement and gravitational acceleration of rigid bodies can be represented with few multivector channels, thanks to the geometric product operation integrated into the architecture. This explains why the architecture can achieve a good performance with very few parameters. However, lowering the loss further requires precise collision detection and modeling. These need substantially more computational operations and a substantial amount of scalar channels, similar to the non-equivariant transformer. This offers a possible explanation for why at a larger compute budget, a model size closer to that of the baseline transformer is compute-optimal.
@@
As much as we would like to, we cannot conclusively settle the question raised in the title of this paper. Our work is limited in several ways. First, we only analyzed a single benchmark problem and two model families. We chose a task with a common symmetry group and general-purpose architectures that are frequently applied to a wide range of problems. We believe it is important to study to what extent our findings generalize to other problems or to other architectures, for instance those based on message-passing over graphs. Moreover, on the problem we studied, we did not set a new state of the art: we deliberately focused on general-purpose models, which do not achieve the same level of performance as highly problem-specific architectures (Allen et al., 2023).
@@
Conclusions
We showed that it can be useful to study strong inductive biases like equivariance through the lens of compute efficiency, rather than only through the established perspective of data efficiency.
Keeping the limitations of our empirical study in mind, we believe that our findings provide evidence that symmetry-aware modeling can be a sensible choice even for large compute and data budgets. To materialize these benefits in practice, it may be both necessary and worthwhile to develop more efficient implementations of equivariant architectures—that is, improve their FLOP throughput.
The benefits and disadvantages of strong inductive biases at scale are important for problems spanning several fields of science and engineering. We hope that our study can encourage further investigations into this question.
@@ Appendix A Dataset
We generate a benchmark dataset of rigid-body interactions with the Kubric simulator (Greff et al., 2022), which is based on the PyBullet physics engine (Coumans & Bai, 2016–2024). We follow the MOVi-B configuration as used by Allen et al. (2023): we generate trajectories trajectories of 96 frames at 48 frames per seconds. Our training set consists of $4\cdot 10^{5}$ such trajectories, while we use 100010001000 trajectories each for the validation and test set. Each trajectory includes between 3 and 10 three-dimensional objects, each consisting of between 98 and 2160 mesh faces. In addition, there is one planar object for the ground floor, canonically placed at $z=0$. The objects are accelerated by collision forces as well as gravity, which canonically points in negative $z$ direction.
@@ Appendix B Models
Relation between equivariant and base model
In many equivariant architectures such as steerable CNNs (Cohen & Welling, 2016), the equivariant linear maps form a subspace of the corresponding space of kernels of a non-equivariant base model. This is not the case for GATr.
However, this difference only holds at the level of linear maps. Nonlinearities generally differ between equivariant and non-equivariant architectures. Therefore, equivariant architectures cannot usually be easily identified with a subspace of the parameter space of a standard non-equivariant architecture.
@@
For the baseline transformer, these features are embedded using random Fourier features (Tancik et al., 2020) with 128 frequencies sampled from a Gaussian with standard deviation 0.10.10.1.
@@
We also experimented with predicting all vertex positions directly, without enforcing object rigidity, as well as with parameterizing elements of the Lie algebra of $\mathrm{E}(3)$, which would then be exponentiated to construct transformations $h$. Both approaches performed worse in initial tests.